OpenStudy (anonymous):

What is the cartesian form of r^2=theta

OpenStudy (ikram002p):

hmm \(\large x^2+y^2=\tan^{-1} \frac{y}{x}\)

OpenStudy (anonymous):

How could i simplify that. ?

OpenStudy (ikram002p):

u dnt need to simplify ( thats why using polar is better in this case) \(\large \tan (x^2+y^2)=\frac{y}{x}\) \(y=\large x\tan (x^2+y^2)\)

OpenStudy (anonymous):

Ohh okay thanks (:

OpenStudy (ikram002p):

got it ?